Identify the t-value given the following. Use the T-table provided in this module.

Confidence Level

Tailed Test

Given

T critical Value

90%

Two - Tailed Test

df = 23

95%

One - Tailed Test

df = 12

99%

Two - Tailed Test

n = 25

90%

One - Tailed Test

n = 11


1
Expert's answer
2022-06-19T16:07:14-0400

a. Based on the information provided, the significance level is "\alpha = 0.10," "df=23" and the critical value for a two-tailed test is "t_c =\pm1.713872."


b. Based on the information provided, the significance level is "\alpha = 0.05," "df=12" and the critical value for an one-tailed test is "t_c =1.782288"

(right-tailed) or  "t_c =-1.782288"(left-tailed).


c. Based on the information provided, the significance level is "\alpha = 0.01," "df=25" and the critical value for a two-tailed test is "t_c =\pm2.787436."


d. Based on the information provided, the significance level is "\alpha = 0.10," "df=11" and the critical value for an one-tailed test is "t_c =1.36343"

(right-tailed) or  "t_c =-1.36343"(left-tailed).